Grade 12Algebra

if x =2 find the value of x+√x+√x+√x........soon.if x =2 find the value of x+√x+√x+√x........soon

Tazim
9 Years agoGrade 12

2 Answers

Khimraj
ApprovedApproved Tutor Answer9 Years ago
Let y = x +√x +√x + √x... Then (y-2)² = x +√x +√x +√x... So y² -4y + 4 = y and y²-5y +4=0.then (y-4)(y-1) =0. And so y = 4. Hope it clears. If u have doubts then please clearify. If u like my answer then plz approve answer.
SAR
5 Years ago
Let Y = x + √x+√x+√x+...     .........1
Given x=2,
 
 We have Y-2=√x+√x+√x+....   .......2
(Eqn.2 is squared on both sides)
Then,(Y-2)2=x+√x+√x+√x+.....
 
From eqn.1 we get
(Y-2)2=Y
Then
Y2-4Y+4-Y = 0,
Then Y2-5Y+4=0
Applying Factorisation theorem 
We get Y=1 or 4.
Hence answer is 1 or 4